OpenStudy (anonymous):

find the tangents of the following curves x^2 + 16y^2 = 32 you get\[y'= \frac{-x}{16y}\]please explain

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

\[2x+16y^2=0\]\[16y^2=-2x\]\[8y^2=-x\]\[y^2=\frac{-x}{8}\]\[y'=\sqrt{\frac{-x}{8}}dx\]

OpenStudy (phi):

use implicit differentiation see http://www.khanacademy.org/math/calculus/differential-calculus/implicit_differentiation/v/implicit-differentiation-1 for \[ x^2 + 16y^2 =32 \\ \frac{d}{dx} \left(x^2 + 16y^2 =32\right) \\ \frac{d}{dx} x^2 + 16 \frac{d}{dx}y^2 = \frac{d}{dx} 32 \\ 2x \frac{dx}{dx} + 16 \cdot 2 y \frac{dy}{dx}= 0 \] that simplifies to the result.